OpenStudy (anonymous):

factor the diffrence of squares? Whats the answer for 25k^2-9 and 16x^2-40x+25

OpenStudy (mathstudent55):

The first one is a difference of squares. The second one is not a difference of squares. For a difference of squares: \(a^2 - b^2 = (a + b)(a - b) \)

OpenStudy (anonymous):

I still dont get it

OpenStudy (mathstudent55):

Look at this term: \(x^2\) Obviously, \(x^2\) is the square of x. A term that is a square can be more complicated. For example, \(4x^2\) is the square of 2x. \(9y^2\) is the square of 3y.

OpenStudy (blackops2luvr):

49?

OpenStudy (anonymous):

for the first one I got 625^2-450k+81

OpenStudy (mathstudent55):

A difference means a subtraction. A difference of squares is a square being subtracted from a square. Look at your first problem, \(25k^2 - 9\) Notice that \(25k^2\) is the square of 5k. 9 is the square of 3. Since 9 (a square) is being subtracted from \(25k^2\) (also a square), you have a subtrtaction of squares, or a difference of two squares.

OpenStudy (mathstudent55):

To factor a difference of squares, follow this pattern: \(a^2 - b^2 = (a + b)(a - b) \) \(25k^2 - 9\) \(= (5k)^2 - (3)^2\) In the line above, you see more clearly that it's the difference of two squares. Now follow the pattern. \( = (5x + 3)(5k - 3) \)